Circular No. 8607

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA SCUTI 2005 M. Soma, National Astronomical Observatory of Japan, reports the

discovery by A. Takao (Kitakyushu, Fukuoka, Japan) of an apparent nova (mag 10.3) on an unfiltered CCD image taken on Sept. 30.522 UT with a 120-mm f/4 telephoto lens, the new object appearing at α = 18^{h}32^{m}.1, δ = −6^{o}43′.6 (equinox 2000.0). Nothing was present at this location on films taken by Takao during the past five years (the most recent of which was taken on Aug. 15; no limiting magnitude given). H. Yamaoka, Kyushu University, writes that the new object appears at mag ∼ 10 on ASAS-3 images taken on Sept. 28.088 and Oct. 1.025, but was not present on Sept. 24.629 (limiting mag perhaps 13); he adds that an independent discovery of the object at mag ∼ 10.9 was made by K. Haseda (Aichi, Japan) from photographs taken with a 120-mm-focal-length f/3.5 lens on Sept. 30.417 and 30.422, the position given as α = 18^{h}32^{m}04^{s}, δ = −6^{o}43′23′′. A. C. Gilmore and P. M. Kilmartin report that short unfiltered CCD exposures taken on Oct. 1.393 with the Mt. John 1-m f/7.7 reflector show the new object at mag 10.0, and they provide the following precise position: α = 18^{h}32^{m}04^{s}.75, δ = −6^{o}43′34′′.3 (from 111 UCAC-2 stars, mean residual 0′′.1 in each coordinate); comparison with a Digitized Sky Survey red plate from 1988 Aug. 10 shows no obvious precursor, though a star of red mag 17.9 lies to the east of the nova (at position end figures 04^{s}.92, 33′′.4). S. Nakano, Sumoto, Japan, forwards the following position end figures for the nova from unfiltered CCD frames taken by K. Kadota (Ageo, Saitama, Japan) on Oct. 1.389 with a 0.25-m f/5.0 reflector: 04^{s}.74, 35′′.0 (mag 9.9).

SUPERNOVA 2005ep IN UGC 10626 Further to IAUC 8605, T. Puckett and M. Peoples report the discovery

of an apparent supernova (mag 18.0) on an unfiltered CCD frame taken with the 0.60-m automated supernova patrol telescope on Sept. 30.05 UT. The new object was confirmed at mag 17.9 on an image taken by D. McDonald, Celbridge, Ireland, with a 0.2-m reflector on Sept. 30.8. SN 2005ep is located at α = 16^{h}56^{m}35^{s}.49, δ = +58^{o}01′30′′.2 (equinox 2000.0), which is 2′′.7 east and 12′′.7 south of the center of UGC 10626. Nothing is present at this location on images taken by Puckett on Sept. 6, 9, 11, and 18 (limiting mag ∼ 20.0).

2005 October 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green